OpenStudy (anonymous):

The students at Midtown Middle school sold flowers as a fundraiser in September and October. In October, they charged $1.50 for each flower. The October price was a 20% increase of the September price. Part A What was the price of the flowers in September?

OpenStudy (anonymous):

@briananicolenelson @angelic._. @jabez177 @natsent06

OpenStudy (anonymous):

@CausticSyndicalist @phi

OpenStudy (anonymous):

Part B The seventh-grade class earned 40% of the selling price of each flower. In September, they sold 900 flowers. In October, they sold 700 flowers. Select a choice from each drop-down menu to make a true statement. The seventh-grade class earned more money in ___ by ___.

OpenStudy (anonymous):

Part B: 1st answer: September, October 2nd answer: 30, 75, 190

OpenStudy (phi):

what did you get for Part A: Part A What was the price of the flowers in September?

OpenStudy (anonymous):

Yes, that's the question

OpenStudy (anonymous):

For part A

OpenStudy (phi):

and what did you get for the answer? do you know what 20% of 1.80 is ?

OpenStudy (anonymous):

There are 2 parts to the question

OpenStudy (anonymous):

no, let me figure it out

OpenStudy (anonymous):

.36?

OpenStudy (phi):

oh, October at 1.50 is the higher price we have to find the Sept price. Call it x: x + 0.2x = 1.50

OpenStudy (phi):

in other words, the sept price x plus 20% of x is october's price 1.50 x+0.2x is the same as 1.2x 1.2x= 1.50 can you find x ? (divide both sides by 1.2)

OpenStudy (anonymous):

x=1.25?

OpenStudy (phi):

yes. Let's check: 20% of 1.25 is 0.2*1.25 = 0.25 so they increased 1.25 by 0.25 to get 1.50 and that matches 1.50 that they told us.

OpenStudy (phi):

in Sept they charged 1.25 for each flower in Oct they charged 1.50 for each flower. now you have to find how much money they got for selling In September, they sold 900 flowers. in Oct, they sold 700

OpenStudy (anonymous):

so then multiply?

OpenStudy (phi):

multiply cost times number of flowers

OpenStudy (anonymous):

September: 1,125 October: 1,050 Correct?

OpenStudy (phi):

yes

OpenStudy (anonymous):

Correct @phi?

OpenStudy (phi):

so it's clear they made more money in Sept. how much more ?

OpenStudy (anonymous):

So then Part A is: $1,125?

OpenStudy (anonymous):

75$ more

OpenStudy (phi):

For part A, I think they want the price per flower, so 1.25

OpenStudy (anonymous):

oh yeah. ok

OpenStudy (anonymous):

and part b

OpenStudy (phi):

for part B, we have to read the question carefully.

OpenStudy (anonymous):

september, 75

OpenStudy (phi):

The seventh-grade class earned 40% of the selling price of each flower.

OpenStudy (anonymous):

its the true statement

OpenStudy (anonymous):

so then the answers are: september, 75

OpenStudy (phi):

they only earned 40% of the numbers

OpenStudy (anonymous):

right???

OpenStudy (anonymous):

oh...

OpenStudy (anonymous):

so then subtract

OpenStudy (anonymous):

so then they made $675

OpenStudy (anonymous):

and in october they made $630 dollars

OpenStudy (phi):

the amount they earned was 0.4 * 1,125 and 0.4 * 1050 the difference is 0.4 * 1,125 -0.4 * 1050= 0.4*(1125-1050)= 0.4*75 = ?

OpenStudy (anonymous):

??

OpenStudy (anonymous):

I don't get it. There's too many numbers

OpenStudy (anonymous):

can you break it down?

OpenStudy (phi):

ok. you have the right idea, but it seems you used 60% ? use 40%

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

but I keep on doing it and I arrive to the same thing

OpenStudy (phi):

you found they sold 900 flowers at 1.25 to get 1125 they earned 40% of that amount, or 0.4*1125 = ?

OpenStudy (anonymous):

450

OpenStudy (anonymous):

and the other one's 420

OpenStudy (phi):

yes

OpenStudy (anonymous):

so then the answer is 30!!

OpenStudy (phi):

yes.

OpenStudy (anonymous):

thx sooooooooooooo much @phi!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (phi):

that line they tossed in about earning 40% was tricky

OpenStudy (anonymous):

yes, true

OpenStudy (anonymous):

could you help with another one?

OpenStudy (anonymous):

ill open another question

OpenStudy (phi):

if you make it a new post. this one is a bit long.

OpenStudy (anonymous):

lol :D

OpenStudy (anonymous):

thx 4 the medal, I gave you one